OpenStudy (anonymous):

Jamal noticed that 3/5 of the children on the playground were boys. Which could be the numbers of boys and girls on the playground? A. 3 boys and 5 girls B. 5 boys and 3 girls C. 8 boys and 5 girls D. 15 boys and 10 girls

OpenStudy (anonymous):

i think A am i right

OpenStudy (timaashorty):

nope it's not A. 3 means the number of boys. 5 means the total of children because It's sayin 3 out of 5 are boys.

OpenStudy (anonymous):

D?

OpenStudy (timaashorty):

this is a tricky question but I do think it's D.

OpenStudy (anonymous):

ok its right but i failed for other questions

OpenStudy (timaashorty):

15+10= 25 15 are boys so 15/25 and then 15/25= 3/5.

OpenStudy (timaashorty):

what other question?

OpenStudy (anonymous):

there were other questions that i did and i failed the test

OpenStudy (anonymous):

this one was right though

OpenStudy (timaashorty):

Oh, goodluck then [:

OpenStudy (anonymous):

thanks